### v1

# Locus: Agentic Predicate Synthesis for Directed Fuzzing

Abstract.
Directed fuzzing aims to find program inputs that lead to specified target program states. It has broad applications, such as debugging system crashes, confirming reported bugs, and generating exploits for potential vulnerabilities. This task is inherently challenging because target states are often deeply nested in the program, while the search space manifested by numerous possible program inputs is prohibitively large. Existing approaches rely on branch distances or manually-specified constraints to guide the search; however, the branches alone are often insufficient to precisely characterize progress toward reaching the target states, while the manually specified constraints are often tailored for specific bug types and thus difficult to generalize to diverse target states and programs.
[…]
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Directed Greybox Fuzzing (DGF) aims to search for program inputs leading its execution to reach specific target program states, e.g., an index variable used to access an array is out of the array bounds, thereby uncovering potential bugs or vulnerabilities. It is widely used in software engineering and security applications, including debugging system crashes(Tan et al., 2023; Wang et al., 2021), testing patches(Böhme et al., 2013; Marinescu and Cadar, 2013), verifying bug reports from static analysis(Christakis et al., 2016; Murali et al., 2024), and generating Proof-of-Concept (PoC) exploits of vulnerabilities(Brumley et al., 2008; Liu et al., 2024a). While the original purpose of directed fuzzing is largely not for discovering new bugs, i.e., it needs a specified target, it has found numerous impactful zero-day vulnerabilities (Liang et al., 2020; Nguyen et al., 2020; Wang et al., 2025; Cao et al., 2023; Shah et al., 2022).
DGF is challenging, as the target program states are often deeply nested in the program, while the search space introduced by the complexity of real-world software is prohibitively large. To speed up the search and schedule promising inputs, most existing techniques leverage the control flow proximity, e.g., the distance to target location in the control flow graph, or heuristics based on the semantics of the branch predicates (Chen and Chen, 2018; Li et al., 2017; Shah et al., 2022), e.g., the target state if x==42 has the distance metric of |x42||x-42|. However, such feedback is sometimes too sparse or indirect to reliably measure the progress, especially when there is a long chain of implicit preconditions guarding the target program states (Mitropoulos et al., 2023; Holler et al., 2012; Dutra et al., 2023; Rigger and Su, 2020; Klees et al., 2018; Zhang et al., 2023; Kim et al., 2023; Sherman and Nagy, 2025). For example, triggering CVE-2018-13785 in libpng requires a PNG file to satisfy a precise sequence of preconditions, i.e., valid signature, correct chunk layout, specific IHDR fields (e.g., bit depth, color type), and a magic image width (0x55555555), to trigger an integer overflow (Kim et al., 2023), while the predicates to explicitly check these conditions are largely absent in the code to provide an incremental progress guidance.
To capture the intricate feedback to improve search efficiency, more advanced approaches identify progress-capturing constraints in the program to drive execution towards satisfying specific temporal orders and preconditions (Meng et al., 2022; Holler et al., 2012; Lee et al., 2021; Liang et al., 2020; Li et al., 2017; Fioraldi et al., 2021; Padhye et al., 2019; Ispoglou et al., 2020). However, these constraints are often manually crafted by experts and tailored to a few specific target state types, e.g., focusing on a temporal memory safety bug like a use-after-free by enforcing an allocate–free–use sequence. As different programs can have diverse target states and disparate functionalities, the feedback metric effective in one case may not generalize to another (Chen et al., 2019).
As Machine Learning (ML) and Large Language Models (LLMs) have demonstrated surprising code reasoning capabilities, there has been a growing interest in extending such capabilities to help guide fuzzing (Patra and Pradel, 2016). A common approach is to employ LLMs to directly generate inputs or grammar-aware input generators (i.e., fuzz driver or harness) (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d; Lyu et al., 2024; Chen et al., 2025; Yang et al., 2025a), where the preconditions useful to reach target states are expressed as part of the input grammar constraints. However, not all conditions to reach target states can be easily represented as input grammars. For example, Figure 1 shows that an intermediate state (!found_plte) necessary to trigger the buffer overflow in libpng only emerges in the middle of the execution and cannot be easily checked at the input level. More importantly, such a task formulation is particularly challenging for LLMs, as reasoning from the target program states all the way to the input often requires an extremely long context and convoluted reasoning chain (Jiang et al., 2024; Ding et al., 2024a; Li et al., 2024a; Risse et al., 2025; Risse and Böhme, 2024), a common pitfall for LLM hallucination (Sahoo et al., 2024; Yao et al., 2025), let alone the challenge to verify the correctness of the LLM generated inputs and harnesses, which can significantly impede the fuzzing progress if the generated constraints are incorrect (Padhye et al., 2019; Ispoglou et al., 2020).
Our approach  We present Locus, a new framework that integrates LLMs’ code reasoning capabilities for directed fuzzing by synthesizing semantically meaningful and verifiable predicates to guide the search. Unlike existing LLM-based approaches that focus on constraining the search space at the input (harness) level, which poses a high reasoning burden on LLMs and can be hard to verify, Locus generalizes the constraint generation to be at arbitrary program points. Specifically, given a target state to reach, Locus automates the reasoning about the intermediate program states that verifiably relax the target states and synthesizes predicates as a curriculum to capture the gradual progress towards reaching them. Such predicates serve as the preconditions dominating all executions to reach the target states, providing fine-grained progress feedback to guide the fuzzers for input scheduling and early termination(Luo et al., 2023; Huang et al., 2022; Srivastava et al., 2022a). As these predicates are implemented as source-level instrumentation, they are agnostic to any fuzzer implementations and can thus be integrated without any customization. Moreover, as the instrumentation is a one-time offline process, the cost of running Locus can be amortized in all the succeeding fuzzing procedures.
Agentic design  Automating the generation of effective progress-capturing predicates for diverse target states involves nontrivial reasoning across the entire codebase, spanning multiple functions and their associated data and control flows. Simple prompting techniques, such as in-context learning, chain-of-thought (CoT) prompting, retrieval-augmented generation, etc., can hardly support sophisticated reasoning at the repository level. To this end, Locus features an agentic synthesizer-validator workflow, equipped with diverse program analysis tools to support the traversal of the control flow graph, tracking data dependencies, retrieving function calls, and symbolic execution. They serve as the agent’s action space, allowing the LLM to reason via CoT and act by calling them (Yao et al., 2023) to iteratively propose, refine, and validate candidate predicates.
Importantly, by constraining the output space of the agent at the predicate-level, Locus ensures any (inevitable) LLM errors are corrected before deployment for fuzzing. Specifically, Locus’s output is validated by both the compiler and symbolic execution (Cadar et al., 2008). The former checks the syntactic correctness, while the latter ensures the generated predicates strictly relax the target states, i.e., by checking whether there exists a path that violates the predicates while satisfying the target states. Such a strict relaxation ensures that the fuzzing execution can be safely early-terminated if it violates the predicate. Figure 2 shows the Locus’s workflow.
[…]
2. Overview
In this section, we first describe the background of directed fuzzing. We then contrast our idea with the existing approaches to demonstrate how Locus complements the existing design (Figure 1).
(a) A simplified CFG
(b) LLM-generated code
(c) Locus synthesizes predicates in relevant functions.
Figure 1. A motivating example (CVE-2013-6954) showing how Locus complements existing works. (a) Traditional approaches based on distance to targets in CFG lack fine-grained guidance to distinguish nodes when they have the same distance. (b) LLM-based harness generation is limited to help reach the target. (c) Predicates (as if statements) synthesized by Locus provide extra semantic guidance for DGFs, while relaxing the constraint generation from input-level to arbitrary program points.
2.1. Directed Fuzzing
Canary  Directed fuzzing aims to generate inputs that drive the program execution to reach predefined program states. In this paper, we rely on the concept canaries (Hazimeh et al., 2020) to explicitly represent the target states, which are considered reached when the corresponding canary condition is satisfied. Early works like AFLGo(Böhme et al., 2017) and Beacon(Huang et al., 2022) treat canaries primarily as the reachability to particular program points, i.e., specific line numbers in a code file. Recent works (Hazimeh et al., 2020; Weissberg et al., 2024) adopt assertion-based canaries, where predicates are explicitly inserted to check vulnerability-triggering conditions at runtime. An example of a canary is highlighted in the bottom-left code snippet of Figure 1.
Canaries for directed fuzzing originate from various sources, including static analysis alerts, manually identified vulnerability sites, or runtime sanitizers. For example, address sanitizers(Serebryany et al., 2012) can also be approximately viewed as canaries checking memory safety violations, e.g., inserting canary(index > maxbound) to detect out-of-bound accesses.
Common strategies to reach canaries  Previous research on directed fuzzing can be broadly grouped into four main strategies:
  • Distance-guided scheduling. This approach approximates the distance from the currently covered regions of the CFG to the canary, and then prioritizes seeds that are more likely to drive execution along paths that reduce this distance(Böhme et al., 2017; Du et al., 2022; Chen et al., 2018). For paths that are unlikely to reach the canary, the execution can be early terminated(Luo et al., 2023; Huang et al., 2022; Srivastava et al., 2022a).
  • Specialized progress-capturing state representations. Some approaches introduce domain-specific abstractions tailored for certain classes of vulnerabilities (Lee et al., 2021; Meng et al., 2022). For example, tracking dataflow from memory allocation to deallocation events for the use-after-free vulnerability.
  • LLM-assisted harness generation. A more recent direction leverages LLMs to synthesize input generators to manipulate raw program inputs (Liu et al., 2024b; Xu et al., 2025; Zhang et al., 2024e; Yang et al., 2025a). These methods attempt to constrain the valid execution at the input level by leveraging the learned input grammar knowledge in the LLMs.
2.2. Motivating Example
We use a real-world vulnerability CVE-2013-6954 in libpng, a widely used C library for parsing PNG files, to demonstrate how Locus complements existing approaches.
A simplified CFG of this vulnerability is shown in Figure 1(a), where all the nodes highlighted in gray are equidistant to the canary node (highlighted in yellow), but only one node PLTE is in the path that can lead to the vulnerability (annotated as green arrows). Based on this, fuzzers can only perform conservative pruning (annotated as dashed arrows) while omitting a large part of the paths that are irrelevant to the vulnerability, e.g., all solid black arrows. Capturing the progress towards reaching this canary requires specialized knowledge about libpng parsing logic. However, developing such specialized progress-capturing state representations requires manual efforts from experts and cannot easily generalize to different vulnerabilities and programs.
Figure 2. Overview of Locus workflow. Locus takes as inputs the program codebase $P$ and the canary $\psi$, and produces a program $P^{\prime}$ instrumented with the progress-capturing predicates. The predicate branches provide extra coverage feedback and guards (via early termination) to guide the fuzzer toward reaching the target state, i.e., canary $\psi$, more efficiently.
Figure 1(b) shows an example of the input generator synthesized by LLM-based approaches that enforced some specific input constraints. However, this approach is inherently limited as program inputs can have sophisticated structures that cannot be easily enforced at the input level, e.g., the PNG file contains a compressed data chunk IDAT. Therefore, the constraints posed on the input generated by the LLM often reduce to generic input grammars, e.g., png_sig_cmp only trivially ensures the input is a valid PNG file. We also note that the synthesized generators cannot be easily checked to determine whether they can effectively help reach the canary. For example, png_write_frame_head constrains the generated input to a special PNG type APNG that is essentially impossible to reach the canary, but automatically checking this fact is challenging. Moreover, to constrain the input to satisfy certain properties necessary to reach the target, the generator sometimes needs to repeat the input processing logic presented in the execution path. Such repeated execution introduces additional overhead.
Locus synthesizes progress-capturing predicates at arbitrary program points to provide more fine-grained guidance and complement the above approaches. Specifically, Locus generates a predicate in png_read_info (shown in Figure 1(c)), representing the precondition to trigger a real-world buffer overflow vulnerability and terminate the execution if it is not satisfied (if(!found_plte)). The target state is highlighted in a canary statement. Locus’s trajectory reveals that it relies on the semantic reasoning of libpng’s parsing behaviors for generating this semantically meaningful predicate.
Specifically, PNG files consist of a series of structured chunks, many of which are optional and can appear in varying orders. One such optional chunk, PLTE, stores the palette data used for indexed-color images and must adhere to strict size constraints. In this example, an input PNG file can trigger a buffer overflow in the png_set_PLTE function only when the PNG file contains a PLTE chunk, which serves as the necessary state before the target state can be reached, i.e., the palette size exceeds the expected bounds.
Existing fuzzers struggle with such a vulnerability due to the complex parsing logic. Since PNG chunks may appear in arbitrary order and many are optional, the primary parsing routine implements a loop in png_read_info to iteratively process each chunk. Within this loop, multiple branches exist, and each is responsible for handling a specific chunk type, e.g., IHDR, IDAT, PLTE. This loop makes it possible to identify whether a given PNG file contains a PLTE chunk. However, since these branches are syntactically parallel and executed without a fixed order, they all appear equidistant from the vulnerability site (png_set_PLTE) in the control flow graph. As a result, the naive path distances cannot distinguish between them to prioritize one path over the other. This explains why, in Table 2 and Table 3, all existing fuzzers take a substantial amount of time to reach this target state in libpng.
Locus starts with generating a predicate to check whether the input PNG file contains a PLTE chunk at the caller of the canary function ( 1). This implies that if the PNG file does not contain a PLTE chunk, the execution can terminate, as it is impossible to reach the target state. While this predicate is semantically correct, i.e., the symbolic execution confirms there is no feasible path to satisfy the target state while violate the generated predicate (Section 3.4), and can help filter out non-palette-based PNG files, it is almost redundant as there is an existing one immediately after the generated check (if (color_type & MASK_PALETTE)), and thus the generated predicate could barely help the fuzzer.
To address this issue, Locus includes another refinement iteration (Algorithm 1) to propagate this predicate to a location closer to the program entry such that the infeasible execution can be terminated earlier. By traversing the call graph and retrieving the functions along the call chain, the refinement proposes a new location in function png_read_info right after the parsing loop ( 2). Such a propagated predicates will also be validated again to ensure it remains a strict precondition to reach the target states. At the next iteration of refinement, Locus proposes to propagate this predicate further inside the branch that parses the IDAT chunk ( 3). That is because the specification of the PNG file requires that the optional PLTE chunk must appear before the IDAT chunk. Therefore, by the time the parsing procedure reaches the IDAT chunk, we can already determine whether the input PNG file contains a PLTE chunk.
[…]
3. Methodology
Table 1. Toolset for synthesizer agent
API Description Output Example Stats
Search
class(cls, [file]) Search for class or structure cls in the codebase or file typedef struct png_XYZ { ... } 13%
method(m, [file]) Search for method or function m in the codebase or file void png_read(png_ptr, ...) {...} 17%
symbol(s, [file]) Search for symbol in the codebase or file TIFFDataType dtype = TIFF_BYTE; 10%
code(c, [file]) Search for code snippet c in the codebase or file raw2tiff.c:302:memcpy(buf, ...) 21%
Graph
callers(f) Get the caller names of function f in the codebase [exif_iif_add_tag, exif_iif_add_fmt, …] 13%
callees(f) Get the callee names of function f in the codebase [php_strnlen, estrndup, php_ifd_double,...] 6%
references(s) Get all references of the symbol s in the codebase [misc.c:2670:if(sig_num!=SIGALRM){, ...] 8%
Listing
files(dir) List all file names under the given path dir [caf.c, chunk.c, common.h, ...] 11%
classes(file) list all classes and structures defined in file [SF_INFO, sf_private_tag, SF_VIRTUAL_IO, ...] 5%
methods(file) List all methods and functions defined in file [exif_get_tag_ht, ifd_get32s, ptr_offset, ...] 6%
Figure 2 illustrates the high-level workflow of Locus. Given a specified canary $\psi$ and the program codebase $P$, Locus outputs a new codebase $P^{\prime}$ instrumented by a set of predicates $\Phi$, where each $\phi\in\Phi$ is represented by a branch condition with early exit if the predicate condition is not satisfied. The synthesizer is responsible for generating candidate predicates $\Phi$, while the validator ensures that $\Phi$ are both syntactically valid and semantically consistent with the $\psi$, i.e., relaxing the canary conditions. The fuzzer will run on $P^{\prime}$ to receive more progress feedback to the canary while enjoying the early termination. In the following, we formalize the task and then elaborate on each design component in Locus.
3.1. Formalization
We use the notation $P\Downarrow_{x}S$ to indicate that the program $P$, when executed with input $x\in X$ sampled from $P$’s input space $X$, can reach a set of program states $S$. Assume triggering a vulnerability $v$ can be characterized as reaching a set of program states $S_{v}$. Given the program under fuzz ($P^{\prime}$) instrumented by $\Phi$, we need to make sure that the early termination introduced in $\Phi$ preserves the same fuzzing behavior on $P^{\prime}$ as that of $P$, i.e., $\Phi$ do not reject any $x\in X$ that would have reached $S_{v}$ in $P$. To this end, we formally define fuzzing admissibility.
Definition 0.
For a vulnerability $v$, the program $P^{\prime}$ is fuzzing admissible to $P$, iff $\forall x\in X,P^{\prime}\Downarrow_{x}S_{v}\implies P\Downarrow_{x}S_{v}$.
While it is unlikely to ensure $P^{\prime}$ is fuzzing admissible to $P$ in general without a pre-defined target vulnerability $v$, we show such a property is well-defined when $v$ is given and explicitly represented by the canary $\psi$.
A program predicate $\phi:s\rightarrow\{\text{True},\text{False}\}$ is a boolean mapping over the program state space. Concretely, any conditions inside the branch statements, e.g., if, while, or assert, can be regarded as predicates, as they evaluate a Boolean expression over the program state. We define a special class of predicates, namely canaries, to characterize vulnerable states:
Definition 0.
A vulnerability canary $\psi$ is a predicate s.t. $\forall s\in S_{v}\Leftrightarrow\psi(s)=\text{True}$.
The goal of directed fuzzing towards $S_{v}$ is equivalent to finding inputs that satisfy the canary $\psi$. To provide semantically meaningful guides to directed fuzzing, we may instrument the program $P$ with an additional predicate $\phi$. Such instrumentation is admissible if and only if $\phi$ is a relaxation of the true vulnerability canary $\psi$:
Definition 0.
A predicate $\phi$ is the relaxation of canary $\psi$, if $\forall s\in S_{v},\psi(s)=\text{True}\implies\phi(s)=\text{True}$.
By definition, $P^{\prime}$ instrumented by $\Phi$ is fuzzing admissible if every predicate $\phi$ is a relaxation of $\psi$. This suggests that fuzzing $P^{\prime}$ is equivalent to fuzzing $P$ while enjoying the additional guidance and early termination introduced by the instrumented predicates.
Theorem 4.
The instrumented program $P^{\prime}$ is fuzzing admissible to $P$, if $P^{\prime}$ is instrumented with $\Phi$, where every $\phi\in\Phi$ is the relaxation of $\psi$.
We next illustrate how our agentic synthesizer-validator workflow can produce an admissible instrumented program $P^{\prime}$ which helps provide rich semantic feedback for the directed fuzzer to reach vulnerable states $S_{v}$.
3.2. Agent Toolset
Existing agentic practices for software analyses, such as bug fixing or fault localization, are often equipped with lightweight command-line tools to perform local reasoning(Paul Gauthier, 2024; Zhang et al., 2024b; Yang et al., 2024b). This is effective because the root cause of a bug is often spatially close to the observable failure, allowing the agent to retrieve relevant context from a narrow portion of the entire codebase. However, in our task formulation, predicates can be synthesized at any program point. As a result, local reasoning is insufficient, and the agent needs additional tools to support analysis of a broader code context.
To equip the agent with long-context code reasoning at arbitrary program points, we provide a suite of tools that Locus can invoke to navigate the entire codebase and retrieve relevant code snippets. Particularly, in addition to common tools like code search and file listing, Locus integrates specialized tools for traversing program graphs, including call graphs and reference graphs. Through the call graph APIs, the synthesizer agent can retrieve function call relationships and reason the interprocedural control flow, while the reference graph API allows it to identify variable usages, pointer dereferences, and data access patterns across the codebase. Table 1 shows the complete toolset and the ratio of their actual usage in our experiments. Among them, graph traversal accounts for over a quarter of all API invocations.
It is important to note that these graphs are only provided as supplementary references to Locus, as they are all derived statically and may miss program behaviors such as dynamic dispatches and indirect calls. We observe that the LLM used in Locus is capable of leveraging its learned knowledge to bridge this gap. For example, in the case of vulnerability TIF002 (see Table 2 and Table 3), Locus successfully resolved the indirect function pointer tif->tif_decoderow to its concrete implementation PixarLogDecode.
Algorithm 1 Scaffold of Locus’s agentic workflow
1:original program $P$, vulnerability canary $\psi$
2:an target-conditional equivalent program $P^{\prime}$
[…]
5:for all $c\in\mathit{C}$ do
6:  $l\leftarrow\textsc{Localize}(c,P)$ $\triangleright$ find the initial program point
7:  $n\leftarrow 0$
8:  repeat
9:   $\phi_{l}\leftarrow\textsc{Generate}(l,c,P)$
10:   while $\neg\textsc{validate}(\phi_{l},\psi,P)$ do $\triangleright$ syntax and semantic
11:     $\phi_{l}\leftarrow\textsc{Generate}(l,\phi,P)$
12:   end while
13:   $l\leftarrow\textsc{Localize}(\phi,l,c,P)$ $\triangleright$ refine, find a better location
14:   $n\leftarrow n+1$
15:  until $l=\text{None}\lor n>\textsc{MaxIterations}$
16:  $\Phi\leftarrow\Phi\cup\{\phi_{l}\}$
17:end for
[…]
3.3. Synthesis
Algorithm 1 elaborates on the workflow of Locus. For each candidate predicate, it iteratively localizes and generates the predicates (lines 4–13), which are then refined and validated for both syntax and semantics (line 8). Once validated, the predicates will instrument the original program to produce a fuzzing-admissible program (line 16).
As we demonstrate in Theorem 4, a predicate $\phi$ should be instrumented at the execution path that can reach the canary $\psi$. To synthesize $\phi$, LLMs must reason about the root cause of the vulnerability and approximate potential execution traces of the program that can trigger it. The execution trace of a program is often long and complex, involving multiple functions and files. Therefore, a naive one-shot synthesis approach may not be sufficient, as the synthesizer only retrieves limited context and proposes primitive predicates, e.g., an initial predicate generated by Locus ( 2 in Figure 1(c)). To address this, Locus employs an iterative localization-generation refinement workflow. In each iteration, the synthesizer generates a predicate that preserves the same semantic meaning while moving its placement closer to the program entry.
Locus first synthesizes an initial set of predicates by analyzing the canary and approximates a list of semantic characteristics of the inputs that are likely related to this canary, e.g., data structures, types, and properties. The synthesizer is then prompted to consider these constraints from multiple dimensions. For example, in Figure 1, besides the predicate, the synthesizer generates multiple constraints, such as requiring that the input PNG file must contain a valid signature. These approximations are progressively concretized and refined as the synthesizer retrieves more relevant code and reasons about the program execution.
For each approximated characteristic, the synthesizer needs to identify an appropriate program point $l$ where the predicate can be expressed in terms of the variables and expressions in scope. Constraining LLM to directly identify the exact program point is often too challenging and unreliable, so the initial stage only asks the synthesizer to select a candidate function rather than a precise program point. Given the reasoning context generated by the LLM so far, the synthesizer attempts to generate an initial predicate $\phi_{l}$ in the candidate function. Once $\phi_{l}$ passes the validation stage (Section 3.4), we refine it in the following iterations. The goal of such a refinement is to identify an earlier program point where the same semantic check can be performed. This allows the program to reject invalid inputs sooner in the execution, enabling the fuzzer to explore more valid mutations within the same time budget.
It is worth noting that with the refinement iteration, a predicate can be all the way refined toward the fuzzing harness. In some cases, Locus can indeed effectively synthesize predicates at the program entry, making it similar in spirit to automated harness generation works such as HGFuzzer (Xu et al., 2025) and InputBlaster (Liu et al., 2024b). However, in most cases, the input constraints cannot be directly accessed at the harness level. For example, in Section 2.2, verifying the presence and content of a PLTE chunk requires parsing internal structures of the input that are only accessible deeper in the program. This necessitates placing predicates at intermediate program points.
3.4. Validation
The validation step is critical to preserving the correctness of the instrumented program and ensuring that the inserted predicate $\phi_{l}$ maintains the fuzzing admissibility. As shown in Algorithm 1 lines 7 to 9, the validator takes as input the candidate predicate $\phi$, the vulnerability canary $\psi$, and the program $P$, and validates whether the predicate is both syntactically and semantically correct. If the predicate passes both checks, Locus will refine the predicate by exploring potentially better program points closer to the program entries (Section 3.3). If it fails, Locus will self-reflect and attempt to regenerate the predicate, using diagnostic feedback collected from the validator.
Syntax validation  The first component of validation is syntactic checking. A predicate that fails to compile cannot be used in a fuzzing campaign, regardless of its intended semantics. To verify this, the program is instrumented by the predicate $\phi$ at the designated program point $l$ and invokes the project’s build system using a predefined command. If the build fails, the associated compiler error messages will be sent to the synthesizer to repair. This diagnostic information typically involves undeclared symbols, type mismatches, or malformed expressions.
Semantic validation  The second component is semantic validation, which confirms that the predicate $\phi$ strictly relaxes $\psi$. It ensures that the predicate does not reject any execution paths that could reach the vulnerability (Theorem 4). As we demonstrated in Definition 3, such validation requires enumerating all possible program inputs, which cannot be done within an acceptable time budget. Therefore, we utilize symbolic execution to find counterexamples of the relaxation.
Theorem 5.
A predicate $\phi$ is not a relaxation of $\phi^{\prime}$, if there exists $P\Downarrow_{x}s$, s.t. $\phi(s)=\text{False}\land\phi^{\prime}(s)=\text{True}$
Specifically, the predicate is not a strict relaxation of the canary if the symbolic execution can find a path that satisfies the negated predicate $\neg\phi$ while the canary $\psi$ still evaluates to true.
It is natural to employ symbolic execution as a formal checker for Theorem 5 to ensure fuzzing admissibility. However, it often incurs substantial overhead by exploring irrelevant execution paths(Cadar et al., 2008), e.g., branches that are unrelated to either the synthesized predicate or the target canary. This excessive path exploration can lead to prohibitively long validation time, incurring additional overhead of deploying Locus. To mitigate this inefficiency, we adopt a strategy inspired by Chopper(Trabish et al., 2018) that skips specified parts of the code and targets only the exploration of paths according to our selection. Specifically, we perform a lightweight reachability analysis on the CFG and prune nodes that are not reachable from either the predicate or the canary. Locus then initiates symbolic execution to explore the path between the $\neg\phi$ and $\psi$.
4. Evaluation
We aim to answer the following research questions:
Table 2. TTE for each vulnerability in the Magma benchmark across different directed fuzzers. T.O. indicates that the fuzzer cannot find the vulnerability within 24 hours. $\varnothing$ indicates that the fuzzer either could not compile the target program or the preprocessing step exceeded 24 hours.
Vul ID AFLGo SelectFuzz Beacon Titan
origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$
PNG003 4 4 1.0 1 3 3 1.0 1 5 5 1.0 1 8 8 1.0 1
PNG006 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
PNG007 72536 9052 8.0 ¡0.01 58770 8537 6.9 0.04 7764 1659 4.7 ¡0.01 17561 11954 1.5 0.04
SND001 T.O. 419 206.5 ¡0.01 7764 5 1725.4 ¡0.01 432 13 33.2 ¡0.01 115 8 14.4 ¡0.01
SND005 1015 205 5.0 ¡0.01 102 4 23.2 ¡0.01 95 16 5.9 ¡0.01 30412 1046 29.1 ¡0.01
SND006 T.O. 3899 22.2 ¡0.01 8519 12 709.9 ¡0.01 T.O. T.O. - - 26171 905 28.9 ¡0.01
SND007 T.O. 3468 24.9 ¡0.01 7785 62 126.4 ¡0.01 T.O. T.O. - - 310 28 11.1 ¡0.01
SND017 4453 1021 4.4 ¡0.01 3235 1518 2.1 ¡0.01 4861 58 83.8 ¡0.01 40955 256 160.0 ¡0.01
SND020 3805 1291 2.9 ¡0.01 694 542 1.3 ¡0.01 5683 41 138.6 ¡0.01 70432 604 116.7 ¡0.01
SND024 T.O. 2577 33.5 ¡0.01 5957 18 332.8 0.05 T.O. T.O. - - 1074 8 127.8 ¡0.01
TIF002 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
TIF005 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
TIF006 80998 17774 4.6 ¡0.01 T.O. T.O. - - 35493 6715 5.3 0.02 66692 65412 1.0 0.04
TIF007 18513 2107 8.8 ¡0.01 1410 413 3.4 ¡0.01 220 76 2.9 ¡0.01 146 109 1.3 ¡0.01
TIF009 32556 16767 1.9 0.04 12067 4144 2.9 ¡0.01 T.O. T.O. - - T.O. T.O. - -
TIF012 47424 18689 2.5 0.02 9118 4787 1.9 ¡0.01 18782 12274 1.5 0.04 2922 2620 1.1 ¡0.01
TIF014 66000 8134 8.1 ¡0.01 58851 7833 7.5 ¡0.01 19035 9672 2.0 0.02 44384 14360 3.1 ¡0.01
LUA004 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - 69594 67442 1.0 ¡0.01
XML001 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
XML003 T.O. T.O. - - T.O. T.O. - - 44056 13295 3.3 ¡0.01 53689 36365 1.5 0.32
XML009 T.O. 77513 1.1 0.6 T.O. 56 1540.1 ¡0.01 32585 6748 4.8 ¡0.01 17942 13879 1.3 0.03
XML017 17 8 2.1 ¡0.01 7 9 0.7 ¡0.01 10 15 0.7 0 23 13 1.8 ¡0.01
SSL001 T.O. T.O. - - 72078 62637 1.2 ¡0.01 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
SSL003 222 231 1.0 0.04 73 69 1.1 0.02 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
PHP004 165 269 0.6 ¡0.01 11 11 1.0 1 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
PHP009 99 203 0.5 0.04 185 153 1.2 ¡0.01 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
PHP011 $\varnothing$ $\varnothing$ - - 26 23 1.1 ¡0.01 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
SQL018 49520 30855 1.6 0.03 53397 6748 7.9 ¡0.01 T.O. T.O. - - 78238 42275 1.9 0.04
[…]
Table 3. TTE for each vulnerability in the Magma benchmark across different coverage-guided fuzzers. T.O. indicates that the fuzzer cannot find the vulnerability within 24 hours. $\varnothing$ indicates that the fuzzer either could not compile the target program or the preprocessing step exceeded 24 hours.
Vul ID AFL++ AFL MOPT Fox
origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$
PNG003 12 8 1.5 0.04 3 3 1.0 1.00 4 4 1.0 1.00 5 5 1.0 1.00
PNG006 106 63 1.7 0.16 T.O. T.O. - - T.O. T.O. - - 6978 1582 4.4 ¡0.01
PNG007 53104 41101 1.3 0.05 55351 18358 3.0 ¡0.01 42811 37522 1.1 0.07 4651 4009 1.2 0.05
SND001 451 19 24.4 ¡0.01 75940 338 224.4 ¡0.01 285 12 23.6 ¡0.01 447 27 16.6 ¡0.01
SND005 1233 193 6.4 ¡0.01 796 9 88.4 ¡0.01 53 9 6.3 ¡0.01 82 15 5.5 ¡0.01
SND006 7026 24 290.3 ¡0.01 T.O. 2251 38.4 ¡0.01 328 17 19.6 ¡0.01 15422 717 21.5 ¡0.01
SND007 810 34 23.6 ¡0.01 T.O. 3163 27.3 ¡0.01 360 16 22.4 ¡0.01 3226 129 25.0 ¡0.01
SND017 841 196 4.3 ¡0.01 3162 833 3.8 ¡0.01 33 20 1.7 ¡0.01 2117 178 11.9 ¡0.01
SND020 640 204 3.1 0.08 2916 1266 2.3 ¡0.01 326 103 3.2 ¡0.01 2068 176 11.8 ¡0.01
SND024 361 18 19.6 ¡0.01 T.O. 2391 36.1 ¡0.01 326 14 23.4 ¡0.01 2642 32 82.6 ¡0.01
TIF002 72896 69963 1.0 0.85 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
TIF005 420 260 1.6 0.04 T.O. T.O. - - T.O. T.O. - - 4744 990 4.8 ¡0.01
TIF006 970 336 2.9 ¡0.01 52926 25366 2.1 0.02 39810 45124 0.9 0.44 22380 7607 2.9 ¡0.01
TIF007 50 26 2.0 0.04 15022 1172 12.8 0.04 155 56 2.8 ¡0.01 8868 290 30.6 ¡0.01
TIF009 17483 5234 3.3 ¡0.01 19302 14098 1.4 0.01 26049 15622 1.7 0.05 7696 3406 2.3 ¡0.01
TIF012 1731 1122 1.5 0.02 28884 16318 1.8 0.09 3639 953 3.8 0.02 1257 637 2.0 ¡0.01
TIF014 2555 682 3.7 ¡0.01 66072 7303 9.0 ¡0.01 1660 788 2.1 ¡0.01 4666 3275 1.4 ¡0.01
LUA004 20046 11939 1.7 0.03 44954 22434 2.0 0.03 12401 2098 5.9 ¡0.01 35998 6937 5.2 ¡0.01
XML001 3720 1330 2.8 ¡0.01 T.O. T.O. - - 37726 16801 2.2 0.03 T.O. T.O. - -
XML003 2373 1394 1.7 0.04 T.O. T.O. - - 56501 18636 3.0 0.03 12252 6844 1.8 0.04
XML009 2668 1452 1.8 ¡0.01 T.O. 22753 3.8 0.06 706 670 1.1 0.65 10827 8766 1.2 0.04
XML017 52 15 3.5 ¡0.01 26 7 3.7 ¡0.01 6 16 0.4 ¡0.01 77 35 2.2 ¡0.01
SSL001 23533 1583 14.9 ¡0.01 81838 T.O. 0.9 - 43952 16033 2.7 ¡0.01 66778 17764 3.8 ¡0.01
SSL003 193 178 1.1 ¡0.01 90 63 1.4 ¡0.01 96 79 1.2 ¡0.01 225 199 1.1 0.33
PHP004 66156 34974 1.9 0.03 82 89 0.9 0.03 760 185 4.1 0.03 $\varnothing$ $\varnothing$ - -
PHP009 11890 5677 2.1 0.02 143 179 0.8 0.21 650 596 1.1 0.91 $\varnothing$ $\varnothing$ - -
PHP011 1626 934 1.7 0.06 108 76 1.4 0.06 19 13 1.5 0.04 $\varnothing$ $\varnothing$ - -
SQL018 10355 4961 2.1 0.04 71038 32073 2.2 0.04 22656 28101 0.8 0.56 15498 8159 1.9 ¡0.01
[…]
  1. RQ1
    Effectiveness: How effective is Locus in accelerating the generation of PoC inputs for given target vulnerabilities?
  2. RQ2
    Cost and performance: What is the time cost and token cost for deploying Locus?
  3. RQ3
    Ablations: How do the individual components of Locus contribute to its overall performance?
  4. RQ4
    Security impact: How can Locus assist in real-world vulnerability detection scenarios?
4.1. Setup
Dataset  We evaluate Locus on the Magma fuzzing benchmark (Hazimeh et al., 2020), which includes a diverse set of real-world vulnerabilities selected from nine popular open-source software projects. We also consider popular libraries in the wild, e.g., libarchive, to evaluate the capabilities of Locus in finding real-world vulnerabilities (see Section 4.6).
Baselines  We select eight fuzzers (see Table 4) as our baseline, covering both the state-of-the-art directed and coverage-guided ones. For each fuzzer, we use the latest stable version available at the time of writing.
For directed fuzzers, we consider AFLGo(Böhme et al., 2017), SelectFuzz(Luo et al., 2023), Beacon(Huang et al., 2022), and Titan(Huang et al., 2024a). We do not include a comprehensive set of directed fuzzers due to the lack of publicly available implementations and the challenges of integrating them into Magma.
We also consider coverage-guided fuzzers, including the widely used AFL(Zalewski, 2020) and AFL++(Fioraldi et al., 2020). To make the evaluation more comprehensive, we include two additional fuzzers. MOPT(Lyu et al., 2019) enhances fuzzing with optimized mutation strategies, and Fox(She et al., 2024) optimizes the seed scheduling as online stochastic control.
Note that Locus alone is not a standalone fuzzer, but focuses exclusively on code transformations for the target software. Therefore, it is agnostic to the fuzzer implementations and can complement any fuzzers (Figure 2).
Table 4. Overview of selected fuzzers in the evaluation
Fuzzer Category Description
AFLGo(Böhme et al., 2017) directed Distance-based seeds scheduling
SelectFuzz(Luo et al., 2023) directed Selective path exploration
Beacon(Huang et al., 2022) directed Fuzzer with efficient path pruning
Titan(Huang et al., 2024a) directed Targets correlations inference
AFL(Zalewski, 2020) coverage-guided Evolutionary mutation strategies
AFL++(Fioraldi et al., 2020) coverage-guided Community-enhanced AFL
MOPT(Lyu et al., 2019) coverage-guided Fuzzer with Swarm Optimization
Fox(She et al., 2024) coverage-guided Online stochastic control
Metrics  We follow the existing directed fuzzing approaches by adopting the Time To Exposure (TTE) to measure the performance of the baseline fuzzers and the improvement introduced by integrating Locus. TTE measures the time taken by a fuzzer to find the input that satisfies the canary condition. To mitigate the inherent randomness introduced by fuzzing, we follow (Hazimeh et al., 2020) by executing 10 independent fuzzing trials per vulnerability sample and report the average TTE. We also employ the Mann-Whitney U Test(Mann and Whitney, 1947) to demonstrate the statistical significance ($p$-value) of the results.
Implementations  We run all the fuzzers with the same initial seed inputs provided in Magma to ensure a fair comparison. Each fuzzing trial is capped at 24 hours, so those that fail to find the triggering input are recorded at this maximum duration. This potentially underestimates the actual TTE for baseline fuzzers, but it offers a lower-bound estimation. Therefore, the actual improvement brought by Locus can be even larger. To eliminate hardware and system-related discrepancies, all experiments are conducted on a dedicated cluster, where each server comes with an Intel Xeon Gold 6126 CPU and 128GB of RAM running Ubuntu 20.04.
We implement Locus using the PydanticAI framework (Colvin, 2025). The synthesizer’s tooling is primarily built on top of Multiplier (Goodman, 2025). We leverage the SVF static analysis framework (Sui and Xue, 2016) to perform a lightweight reachability analysis and use KLEE (Cadar et al., 2008) as the symbolic execution engine to perform semantic validation. We set the LLM generation temperature to zero to avoid non-determinism. We use o3-mini-2025-01-31 as the default model with reasoning level to medium, but we also evaluate other LLMs in Section 4.4.
4.2. RQ1: Vulnerabilities Reproduction
We apply Locus on both directed fuzzers and coverage-guided fuzzers, and compare the performance w/ and w/o Locus. The results are shown in Table 2 and Table 3.
[…]
4.3. RQ2: Cost Analysis
We measure the time cost introduced by Locus and the token cost incurred by LLM inference.
Time cost  In Section 4.2, we focused on measuring TTE, i.e., the time required to trigger the target canary during fuzzing. However, only focusing on the fuzzing phase can be misleading when evaluating the overall effectiveness of a directed fuzzer. Many directed fuzzers perform expensive static analysis on the target program before starting fuzzing (Böhme et al., 2017; Huang et al., 2024a, b). Likewise, Locus requires additional preprocessing steps, including codebase indexing, symbolic validation of predicates, and agentic predicate synthesis. While such preprocessing costs are often arguably a one-time effort that can be amortized during fuzzing, in practice, we observe that some approaches incur excessive analysis time, sometimes even longer than the time required to discover the bug by the fuzzer.
[…]
LLM token cost  We assess the feasibility of deploying Locus in practice from a financial perspective, focusing on the monetary cost of LLM token usage. Table 5 shows the token cost for our Locus workflow. Locus takes 457k tokens (equivalent to \$0.72 USD) to generate predicates for one sample in the Magma benchmark on average. The monetary token costs show the potential for Locus to act as an affordable step in assisting fuzzing.
Table 5. Average deploy cost for Locus and directed fuzzers. All times are calculated in seconds. T.O. indicates that the fuzzer failed to instrument the target library within 24 hours.
Target PNG SND TIF LUA XML SSL PHP SQL
Size (LoC) 95k 83k 95k 21k 320k 630k 1.6M 387k
Index 11 34 82 9 76 146 244 137
Synthesis 373 331 212 178 215 384 412 349
Validation 261 133 231 280 475 824 353 407
Total 645 498 525 467 766 1354 1009 893
#Tokens (k) 309 303 256 176 653 598 894 467
AFLGo 122 673 2689 85 5608 24799 T.O. 15630
SelectFuzz 84 199 1167 44 807 2597 4554 383
Beacon 64 113 171 35 1656 T.O. T.O. 3721
Titan 96 186 967 49 2936 T.O. T.O. 4965
4.4. RQ3: Ablations
We ablate the design choice of different modules in Locus and study how they generalize to various LLMs in Table 6. We pick five representative vulnerabilities that fall under different vulnerability categories in Common Weakness Enumerations (CWEs). The evaluation compares two representative mainstream fuzzing tools, AFL++ from the coverage-guided fuzzers and SelectFuzz from the directed fuzzers.
Table 6. TTEs by ablating different designs and models.
PNG007 SND001 TIF012 TIF014 SQL018
Abalate different designs
AFL++
origin 53104 451 1731 2555 10355
+base 54830 389 1904 1226 11573
+refine T.O. 23 4417 33931 54268
+valid 41101 19 1122 682 4961
SelectFuzz
origin 72078 7764 9118 58851 53397
+base 79348 2640 8938 12630 48059
+refine T.O. 6 10805 T.O. T.O.
+valid 8537 5 4787 7833 6748
Abalate different LLMs
AFL++
origin 53104 451 1731 2555 10355
w/ o3-mini 41101 19 1122 682 4961
w/ DR1 45104 53 992 1439 5433
w/ Gemini 38215 130 1517 2454 5274
SelectFuzz
origin 72078 7764 9118 58851 53397
w/ O3-mini 8537 5 4787 7833 6748
w/ R1 23699 6 3562 3518 7023
w/ Gemini 7020 12 7206 40332 3921
Effectiveness of each design component  We ablate the individual design and measure the average TTE obtained by the resulting predicates. We begin with the base setting, where we only keep the initial synthesized predicate (Algorithm 1 line 7). Next, we apply the refinement strategy, allowing the synthesized predicate to propagate to a better program point without validating its semantic correctness. Finally, we evaluate the full pipeline of Locus including both the syntax and semantic validation.
Table 6 shows that the design choices in Locus consistently improves the performance. Specifically, we observe that while predicates generated under the base setting may occasionally accelerate fuzzing, they are not consistently beneficial. Without validation, Locus sometimes generates false predicates to the program, as evidenced by the cases PNG007, TIF014, and SQL018.
Varying models  Besides the default o3-mini model, we consider Deepseek R1(DeepSeek-AI et al., 2025) and Gemini 2.0 Flash(Google DeepMind, 2024) to study the generality of our agentic framework to different LLM architectures. Table 6 shows that Locus generalizes to different LLM architectures, except for one case where Gemini fails to bring significant improvement to TIF014. We investigate this case and find that Gemini failed to elevate the generated predicate to the caller closer to the input, such that the additional overhead introduced by evaluating the predicates outweighs the benefit it brings to the fuzzing.
4.5. RQ4: Detecting New Vulnerabilities
We integrate Locus into a real-world vulnerability detection workflow and apply the pipeline to a set of well-fuzzed targets, such as VLC(VideoLAN, 2025), libarchive(libarchive contributors, 2025), and libming(libming contributors, 2025). To construct meaningful fuzzing targets, we generate new canaries through two primary approaches. First, we leverage alerts produced by the static analysis tool SVF(Sui and Xue, 2016), which identify potentially vulnerable program points based on memory access patterns, aliasing behavior, or use-after-free risks. Second, we derive reachability canaries for program points associated with previously patched bugs. The rationale behind this strategy is that many real-world bugs occur in clusters or evolve from incomplete fixes(Wang et al., 2020).
We launch 42 fuzzing samples in total, with each fuzzing campaign lasting for 24 hours. We found seven previously undiscovered bugs, including memory leaks, use-after-free, and out-of-bound access (Table 7). At the time of writing, we responsibly reported all the new vulnerabilities to the relevant maintainers, and one of the maintainers has already drafted a fix patch. To further evaluate the efficiency improvement, we also repeat the fuzzing campaign by fuzzers without Locus. The results show that without Locus, AFL++ alone can only detect 2/7 vulnerabilities, and SelectFuzz alone can only detect 3/7 vulnerabilities. Section 4.6 elaborates on one newly detected vulnerability.
Table 7. New vulnerabilities detected fuzzers with Locus. For each newly found vulnerability, we run AFL++ and SelectFuzz w/ and w/o Locus. We exclude the detailed software information per vulnerability report guidelines.
ID Vul. Type AFL++ SelectFuzz
origin Locus origin Locus
VLC-29163 Memory leak T.O. 30847 T.O. 26317
VLC-29162 OOB access T.O. 74835 T.O. T.O.
VLC-29238 Memory leak T.O. 21085 53872 24766
VLC-29239 Use-after-free 43680 3946 22983 5405
libming-365 Null deref T.O. 80241 T.O. T.O.
libarchive-hvqg Null deref 83622 34327 62748 18309
libarchive-fm54 OOB access T.O. 16397 46577 23280
4.6. Case Study
[…]
Upon further investigation, we find that their employed termination mechanism can sometimes conflict with the canaries. Specifically, Beacon and Titan perform static analysis to identify functions potentially reachable from the vulnerability canary, and primarily prioritize program inputs that can reach the statically identified functions. The predicates synthesized by Locus thus offer only limited guidance when they are put in the functions that are not statically identified as close to the canary.
Figure 3. Locus can sometimes synthesize canary conditions even more precisely than those in Magma, even when provided only with the security patch. Prior to the fix, a dequoting flag was erroneously set to 1, enabling attackers to access uninitialized memory by forcing dequoting.
Canary generation by LocusSection 4.5 discusses our attempts to generate canaries for arbitrary types of vulnerability definitions without relying on existing ones. This shows that Locus does not rely on a pre-defined canary to be applicable. Instead, it can be extended to automatically generate canaries based on diverse representations of the target states, such as the patch and bug descriptions.
To apply Locus in such settings, we also consider letting Locus generate canary conditions based on available security patches. Specifically, we iterate over all security patches in Magma and prompt Locus to automate the generation of a canary condition (where Magma relied on manual analysis), such that the unpatched version satisfies the generated canary, while the patched version violates it. By manually checking all the canaries generated by the LLMs, we confirm that 27 out of 28 security patches can be correctly translated into the same condition.
We find the remaining case to be intriguing: the LLM produces a more precise vulnerability canary than the original one in the Magma benchmark, as shown in Figure 3). Specifically, the vulnerability is found in SQLite3 and classified as CWE-908, i.e., use of the uninitialized resource. When building a collate token, a constant value 1 is incorrectly passed in instead of the dequote flag. This allows attackers to access uninitialized memory via forcibly dequoting, which would lead to memory corruption. The original vulnerability canary only checks whether dequoting is disabled, and invalidates cases for unquoted names since they would not be dequoted even if the dequote flag were set.
Instead, the canary generated by Locus focuses specifically on cases in which a quoted string is mishandled. Such a canary characterizes the real vulnerability more precisely and can potentially reduce the false positives introduced by the imprecise ground truth canaries.
Real-world vulnerability  We elaborate on the vulnerability libarchive-fm54 in Table 7, an out-of-bounds error can occur when the program processes a specially crafted RAR file. The vulnerability stems from insufficient bound checking in the delta filter logic used during decompression. In the RAR format, filters are lightweight data transformation routines applied to improve compression efficiency. One common example is the delta filter, which is often used on binary data, such as audio or image streams, to convert absolute values into relative differences, thus making patterns more compressible. During decompression, the filter operates on two memory regions: the source buffer src, which holds the raw decompressed data, and the destination buffer dst, which stores the filtered output. This vulnerability will be triggered when src inadvertently points into the dst buffer. This overlap can lead to undefined behavior, including memory corruption, out-of-bounds access, or program crashes, as the filter may read from or write to memory locations outside the valid range.
Figure 4 shows the predicates synthesized by Locus that check the input is a valid RAR format file and that the delta filter processing logic will be triggered during decompression. It is seemingly more helpful to generate the predicate to compare the address of src and dst. However, these pointers are only assigned and computed within the vulnerable function itself. By the time this function is reached, it can be too late for such a predicate to effectively guide the fuzzer.
Figure 4. A previously unknown vulnerability in libarchive
5. Related Work
Direct grey-box fuzzing  Directed grey-box fuzzing is particularly challenging, primarily due to the prohibitively large search space with sparse rewards(Li et al., 2024b; Luo et al., 2023), i.e., unlike undirected fuzzing where any new coverage indicates progress. Most prior works develop heuristics through static analysis by computing path distances to target states to narrow down the search space (She et al., 2024; Shah et al., 2022; Huang et al., 2024a; Böhme et al., 2017; Luo et al., 2023; Srivastava et al., 2022b; Li et al., 2024b; Chen and Chen, 2018; Lee et al., 2021). For example, AFLGo(Böhme et al., 2017) introduces the idea of using distance metrics between test inputs and target basic blocks to prioritize seeds that are closer to the target, while CAFL(Lee et al., 2021) further improves this metric by introducing specialized progress-capturing state representation and calculating the distance from the testing inputs to the nearest state.
Similar to Locus, some recent approaches(Meng et al., 2022; Peng et al., 2018; Ba et al., 2022) also explore rewriting the program to direct execution towards hard-to-reach states. These approaches focus on checking sophisticated conditions uncovered by human experts, e.g., temporal properties in network protocols, and may not generalize to broader types of target states and can also incur expensive manual effort. In contrast, Locus offers a generalizable framework for diverse types of programs and target states with minimal human intervention.
LLM-based fuzzing  LLMs have demonstrated promising code reasoning capabilities(Ding et al., 2024c; Wang et al., 2024; Ding et al., 2024b; Gong et al., 2024; Gu et al., 2024; Yang et al., 2025b). There has been growing interest in applying them to support fuzzing (DARPA, 2024). Most existing approaches utilize LLMs to directly generate test inputs for the target program (Zhang et al., 2024c; Hu et al., 2023; Zhang et al., 2024a; Yang et al., 2023, 2024a; Deng et al., 2024, 2023; Xia et al., 2024) or build grammar-aware input generators (i.e., fuzzing harness) to constrain the search space (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d). While these methods demonstrate the potential of LLMs to accelerate fuzzing, they often require LLMs to reason about the target program states all the way to the inputs, making them susceptible to fundamental challenges in program analysis, such as path explosions and alias dependencies (Jiang et al., 2024; Ding et al., 2024a; Li et al., 2024a; Risse and Böhme, 2024), while also pose challenges to LLMs in terms of the bloated context length and unbounded errors.
As opposed to solely generating inputs or fuzzing harnesses, Locus advances LLM-based fuzzing by constraining its generation at the level of progress-capturing predicates. Such a task formulation enables the LLM to operate within the (relatively) local context, allowing Locus to detect certain code behaviors that only emerge midway through execution to inform subsequent input searches, and also facilitates the rigorous validation of potential LLM errors. As a result, this strategy can effectively sidestep reasoning across lengthy function call chains, enabling a greater focus on more local, intermediate function contexts.
LLM-driven proof synthesisLocus shares a similar spirit with recent research that focuses on reasoning about loop invariants and program specifications to synthesize proofs and automate program verification(Chakraborty et al., 2023; Chen et al., 2024; Lu et al., 2024; Kozyrev et al., 2024; Yang et al., 2024c). Locus resembles these approaches in the sense that they also integrate LLMs with rule-based verifiers, such as theorem provers or SMT solvers, to check the validity of synthesized properties. However, Locus relaxes the requirement that the synthesized specifications must facilitate rigorous proof steps, but treats the predicates as best-effort guidance for input search. It bridges the gap between formal verification and dynamic testing, leveraging the reasoning capabilities of LLMs to identify meaningful program states while prioritizing efficiency and generality over formal guarantees.
6. Threats to Validity
Threats to predicate validation  While Locus leverages symbolic execution to bound the error of the synthesizer, it inherits all of its limitations. For example, we still face the path explosion problem, especially when there are loops between the generated predicates and the canary. While we adopt the common strategies to address these limitations, e.g., loop unrolling, we cannot formally guarantee that the relaxation brought by the generated predicates is always valid and can thus effectively guide fuzzing. This leads to orthogonal but interesting future directions on loop invariant reasoning and function summary (using LLMs) for symbolic execution.
Threats to LLM types  Second, our evaluation is currently limited to a specific set of LLMs and benchmark programs. The generality of our results to other models, especially open-weight LLMs and broader software systems, remains to be validated. The performance of our approach may depend on the underlying LLM capability and training data. Additionally, our evaluated benchmark Magma, though built on popular software and real-world vulnerabilities, may not capture the comprehensive challenges present in diverse software.
Threats to data leakage  While the task formulation, i.e., predicate synthesis for directed fuzzing, is arguably hard to suffer from the data leakage problem, the software project in our evaluation is largely included in the training data of any major LLMs. As the ultimate goal is to detect and confirm vulnerabilities, we believe that such a threat is minor compared to the practical security impact, as we have shown in Section 4.6.
7. Conclusion
This paper presented Locus, a novel framework for enhancing directed fuzzing through synthetic progress-capturing predicates. With an agentic synthesizer-validator architecture, Locus effectively guides fuzzing via the predicates while also ensuring any errors in the synthetic predicates are fixed by symbolic execution. Our evaluation demonstrated that Locus substantially improves the state-of-the-art fuzzers. So far, it has uncovered seven previously unpatched vulnerabilities across three software, with one already acknowledged with a draft patch.